OpenStudy (anonymous):

Can someone help me find the solution to sqrt(2x+3) = sqrt(x)?

OpenStudy (anonymous):

\[\sqrt{2x+3}=\sqrt{x}\]

OpenStudy (mrnood):

square both sides of the equation

OpenStudy (anonymous):

okay, so now I should have 2x+3=x right?

OpenStudy (anonymous):

By inspection, x = -3

OpenStudy (mrnood):

so now move all terms with x to one side, and all terms without to the other (be careful with signs...)

OpenStudy (anonymous):

so -x+2x+3=0

OpenStudy (anonymous):

?

OpenStudy (mrnood):

so now move all terms with x to one side, and all terms without to the other (be careful with signs...) (you only did half of that before.... :-)

OpenStudy (anonymous):

-x+2x=-3 ??

OpenStudy (mrnood):

yes - now simplify the left side..... and you're there

OpenStudy (anonymous):

I think x=-3, yes?

OpenStudy (mrnood):

ok - well done

OpenStudy (anonymous):

:) Thanks